# Gravitational radiation from hyperbolic encounters in the presence of dark matter

### v1

@@ ABSTRACT
In this study, we look into binaries undergoing gravitational radiation during a hyperbolic passage. Such hyperbolic events can be a credible source of gravitational waves in future detectors. We systematically calculate fluxes of gravitational radiation from such events in the presence of dark matter, also considering the effects of dynamical friction. We also investigate the binary dynamics through the changes in the orbital parameters by treating the potential due to dark matter spike and the dynamical friction effects as a perturbation term. An insight into the effects of such a medium on the binaries from the corresponding osculating elements opens up avenues to study binary dynamics for such events.
@@ 1 Introduction
Our universe is not empty but filled with non-trivial matter medium. Dark Matter (DM) is the most ubiquitous, whose nature has yet to be fully known, and the physics is still speculative. One of the most sought-after avenues deals with modelling these dark matter particles as WIMPs (Weakly Interacting Massive Particles) and exploring their detection prospects both directly and indirectly[53, 54, 45, 46, 47, 48, 49, 50, 51, 52]. The density profile of such WIMPs is universal, with a cusp near the galactic centre and also forming a spike near the central massive black hole. In recent times the advent of data from the LIGO-VIRGO detectors has been much interest in probing DM with GW [55, 56]. Studies in [57] also supported the fact that phases of GWs can be modified, and such modifications can impact future space-borne GW experiments. In [58, 59], dynamical friction effects from such dark matter spikes showed changes to the phase of a GW from binaries with circular orbits.
@@
We organize the paper as follows: In Sec. (2), we briefly outline the basic setup for such binaries moving in hyperbolic tracks and compute the radiation fluxes from such systems. In Sec. (3), we briefly discuss the particular model of the DM distribution that we will consider throughout this paper. In Sec. (4), we compute GW fluxes due to hyperbolic encounters by taking into effect of the surrounding DM medium. Then in Sec. (5), we include the effect of this DM medium (as well as the GW backreaction) as a perturbative term in the Keplerian orbit equations and calculate the changes in orbital parameters due to these extra effects. We also discuss the details of the numerical method we used to solve the simultaneous equations and list our conclusions therein. Finally, we summarize our results and conclude with future directions with Sec. (6).
[…]
@@ 3 Model for the dynamical friction due to DM spike
We begin this section by exploring the dynamics of binaries moving in a hyperbolic orbit in the presence of a medium, more precisely focusing on the fluxes of radiation. A natural question to ask is about the nature of the medium. Around $27\%$ of the universe’s total mass seems to come from the dark matter, and there are observational evidence in favor of it[63, 64, 65, 66, 67, 69, 68, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. They are ubiquitous, and more interestingly, Black Holes (BHs) in such DM environments build up halos around them due to the effect of the strong gravity. For our studies in this paper we will consider Navarro-Frenkel-White (NFW) profile [87] which form spikes due to the presence of certain overdense regions[88, 89, 90, 91, 92]. These overdense regions, in turn, result from the adiabatic growth of supermassive BHs with masses in the range $10^{6}\sim 10^{9}M_{\odot}$ [88]. It seems logical that an object moving around a central BH might have its orbital motion affected by such spike structures. Optical observations of such activities can lead to indirect tests of the existence of such DM spikes and constrain the spike’s density profile [63, 93, 94, 95]. LIGO/VIRGO and other high-end detectors are gearing up to detect GWs from such binaries[96, 97, 57, 98, 99, 100, 101, 102, 103, 104, 106, 107, 109, 105, 108], shedding light on DM features based on GW observations.
@@
[…]
@@ 5.1 Effect of gravitational potential due to the Dark Matter Minispike
The above equation (5.6) tells us the perturbing force spans the plane in which the binaries are moving and the components contributing to the perturbing force can be easily read off by comparing with (5.2) and identifying $\hat{n}$ with $\hat{r}.$ Then from (5.4) and (5.5) we get,
@@ 5.2 Dynamical friction and accretion effects
Given a binary moving through a medium (which for us is a dark matter medium) it will experience a drag or dynamical friction force. Having discussed the effects of dark matter as perturbing force, we can now focus on the effects of such matter medium on the overall dynamics in such hyperbolic tracks. To this end as we discussed in Section 3 the ”frictional” force of the medium can be modeled after a force law which takes the form -
@@
Once we have these components, we can plug them in the osculating equations (5.3) and (5.4) and find the change in the orbital parameters:
@@ 5.4 Results: Numerical Analysis and Inferences
Now that we have all the relevant equations for studying the binary dynamics, specifically the effects of these various forces due to the presence of dark matter on the rate of change of the eccentricity. Before discussing our result, please note that we have used the fourth-order Runge-Kutta method to solve the coupled differential equations for each case separately, as mentioned in (5.7), (5.11) and (5.14). For each case, three coupled differential equations govern the evolution of eccentricity $e$, impact parameter $b$ and angle of closest approach $\phi_{0}$ w.r.t the true anomaly $\tilde{\phi}.$ Since we have taken true anomaly as our independent variable instead of time, all the plots are with w.r.t $\tilde{\phi}.$ The range of the true anomaly is from $\tilde{\phi}=-\phi_{0}$ to $\tilde{\phi}=\phi_{0}$ indicating the fact that the second object comes from infinity and means binaries after the encounter flies to infinity again. For the net effect, we first add the right-hand side of the (5.7), (5.11) and (5.14) and then sole the resulting coupled differential equation using the fourth-order Runge-Kutta method.

### v2

@@ ABSTRACT
In this study, we look into binaries undergoing gravitational radiation during a hyperbolic passage. Such hyperbolic events can be a credible source of gravitational waves in future detectors. We systematically calculate fluxes of gravitational radiation from such events in the presence of dark matter with different profiles, also considering the effects of dynamical friction. We provide an estimate for the braking index and show how it evolves due to the presence of the dark matter medium. We also investigate the binary dynamics through the changes in the orbital parameters by treating the potential due to dark matter spike and the dynamical friction effects as a perturbation term. An insight into the effects of such a medium on the binaries from the corresponding osculating elements opens up avenues to study binary dynamics for such events.
@@ 1 Introduction
Our universe is not empty but filled with non-trivial matter medium. Dark Matter (DM) is the most ubiquitous, whose nature has yet to be fully known, and the physics is still speculative. One of the most sought-after avenues deals with modelling these dark matter particles as WIMPs (Weakly Interacting Massive Particles) and exploring their detection prospects both directly and indirectly[53, 54, 45, 46, 47, 48, 49, 50, 51, 52]. The density profile of such WIMPs is universal, with a cusp near the galactic centre and also forming a spike near the central massive black hole. In recent times the advent of data from the LIGO-VIRGO detectors has been much interest in probing DM with GW [55, 56, 57]. Studies in [58] also supported the fact that phases of GWs can be modified, and such modifications can impact future space-borne GW experiments. In [59, 60], dynamical friction effects from such dark matter spikes showed changes to the phase of a GW from binaries with circular orbits.
@@
We organize the paper as follows: In Sec. (2), we briefly outline the basic setup for such binaries moving in hyperbolic tracks and compute the radiation fluxes from such systems. In Sec. (3), we briefly discuss the particular model of the DM distribution that we will consider throughout this paper. In Sec. (4), we compute GW fluxes due to hyperbolic encounters by taking into effect of the surrounding DM medium. In Sec. (5), we provide an estimate of the braking index, which may provide a possible observational signature of dark matter (through its dissipative effects) via GW wave signal emitted from hyperbolic encounters. Finally, in Sec. (6), we include the effect of this DM medium (as well as the GW backreaction) as a perturbative term in the Keplerian orbit equations and calculate the changes in orbital parameters due to these extra effects. We also discuss the details of the numerical method we used to solve the simultaneous equations and list our conclusions therein. Finally, we summarize our results and conclude with future directions with Sec. (7).
Notation and Conventions: We use units where the Gravitational constant and Planck constant is set to unity: $\textit{G}=\hbar=1.$ Also, the solar mass value $M_{\odot}$, that we have used all over the text is taken to be $10^{30}$ kg.
@@ 3 Model for the dynamical friction due to DM spike
We begin this section by exploring the dynamics of binaries moving in a hyperbolic orbit in the presence of a medium, more precisely focusing on the fluxes of radiation. A natural question to ask is about the nature of the medium. Around $27\%$ of the universe’s total mass seems to come from the dark matter, and there are observational evidence in favor of it[64, 65, 66, 67, 68, 70, 69, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. They are ubiquitous, and more interestingly, Black Holes (BHs) in such DM environments build up halos around them due to the effect of the strong gravity. For our studies in this paper we will consider Navarro-Frenk-White (NFW) profile [88] which form spikes due to the presence of certain overdense regions[89, 90, 91, 92, 93]. These overdense regions, in turn, result from the adiabatic growth of supermassive BHs with masses in the range $10^{6}\sim 10^{9}M_{\odot}$ [89]. It seems logical that an object moving around a central BH might have its orbital motion affected by such spike structures. Optical observations of such activities can lead to indirect tests of the existence of such DM spikes and constrain the spike’s density profile [64, 94, 95, 96]. LIGO/VIRGO and other high-end detectors are gearing up to detect GWs from such binaries [164, 163, 58, 102, 109, 110, 111, 112, 113, 114, 116, 117, 119, 115, 118, 165], shedding light on DM features based on GW observations.
@@
where $\rho_{sp}$ is a normalization constant, and $r_{sp}$ is the radius of the mini spike. There are estimates as to what this $\alpha$ value can be. Typically this range is: $2.25\leq\alpha\leq 2.5$ [89, 91, 124, 125]. Henceforth we will strictly follow this range. For our subsequent analysis we take the the mass of the central massive object as $m_{1}=10^{3}M_{\odot}$ and the mass of the secondary as $m_{2}=10M_{\odot}.$ Also, $\rho_{sp}=226M_{\odot}/pc^{3}$ and $r_{sp}$ to be around $0.54\,pc$ following [58, 102, 115].
@@ 4 Calculation of Fluxes in the presence of dark matter
5 Possible Observable signatures in GW signal
Having calculated the necessary fluxes associated with the GW emission in the presence of dark matter, including the effect of dynamical friction, we are in a good position to explore the possibilities of detecting such effects in the GW signal. In the context of pulsars, the so-called braking index $n_{b}$, has been a useful quantity to infer about the energy loss mechanism and these mechanisms are mainly mediated by GW emissions. Since pulsars can also spin-down through GW emission associated with asymmetric deformations, it is appropriate to take into account this mechanism in a model that aims to explain the measured braking indices [143, 144]. The values of these braking indices can be $3.15\pm 0.03$, as reported in [145] for PSR J1640-46301, which was surprisingly the largest among the only eight of the $\sim$ 2400 known pulsars measured. This is expected from the pure magneto-dipole radiation model reported in [146, 147, 148, 149, 150, 151, 152]. There have been many interpretations of this, such as those that propose that it may be due to the accretion of fallback material via a circumstellar disk [153], or via the quantum vacuum friction effect [154] or perhaps relativistic particle winds [155, 156] or modified canonical models [157, 158]. Studies in [159, 160] have also indicated that it might be due to the effect of the change in the magnetic moment with time, either through a change in the surface field strength or the angle between the magnetic and spin axes. Taking a cue from this, [161] showed that there could be possible combinations of gravitational and electromagnetic contributions to spindown, which could explain the measured braking indices.
Once we have become familiar with the role of the braking index in inferring the characteristics of binaries, we can explore this in our context. Given the binaries (and the binaries are moving on a hyperbolic track), one can calculate the braking index using the following formula:
$n_{b}=\frac{\mathcal{F}\ddot{\mathcal{F}}}{\dot{\mathcal{F}}^{2}}$ (5.1)
where $\mathcal{F}$ is the orbital frequency. In the context of elliptic orbits, $\mathcal{F}$ is related to the semi-major axis: $a$, via the relation $\mathcal{F}=\frac{1}{2\pi}\sqrt{\frac{M}{a^{3}}}\,.$ It is tricky to define the frequency for a hyperbolic orbit. The period is essentially infinite if a particle comes from infinity and goes to infinity after some scattering process following a hyperbolic track. However, there are many ways one can make a segue way out of this, like modeling the compact binaries after a semi-periodic source[162] or maybe focusing on a finite part of the hyperbolic track like going from point $x$ to $y$ and calculating the time taken for such a path. For a hyperbolic orbit, one can still redefine the usual orbital parameters like mean motion $n$ and the mean anomaly $\mathcal{M}$ as $\mathcal{M}=nt$ and $n=\sqrt{\frac{M}{-a^{3}}}$ which in turn will give us $\mathcal{F}=\frac{1}{2\pi}\sqrt{\frac{M}{-a^{3}}}$ and the period along the finite part of the track being $T=\frac{2\pi}{n}$ We also choose:
$r=\frac{a(e^{2}-1)}{1+e\cos\phi}$ (5.2)
with $e>1$ for the hyperbolic case. Furthermore, the energy of the orbiting particle in these tracks is expressed as $\frac{\mu M}{2|a|}$ ($>0$, as expected for an unbound orbit). Given these, we can express the braking index in terms of $a$ in the following way:
$n_{b}=\frac{5}{3}-\frac{2}{3}\frac{a\ddot{a}}{\dot{a}^{2}}\,.$ (5.3)
The next step towards computing the braking index is to find $a,\dot{a},\ddot{a}$, and the method is straightforward. Assuming that while losing energy through radiation, we neglect the osculations of the orbit and the orbital timescale is such that we can use the Keplerian equations while taking the orbital average, we can calculate the dissipative forces acting on the binaries. We can model these dissipative forces by making them a function of $r$ and $v$ and then taking the average over the hyperbolic orbit. The energy and angular momentum corresponding to these dissipative forces can be calculated as follows [^7] Later, we will convert the $t$ integral into a integral of $\phi$ using chain-rule, with the limits of integration changing from $t=(0,T)$ to $\phi=(0,2\phi_{0})$, where $\phi_{0}$ is shown in Fig. 1 and the corresponding expression is given in (2.3). One can use the expression given in (LABEL:eq:2.1) to replace $\dot{\phi}$ in the integral.:
$\Big{\langle}\frac{dE}{dt}\Big{\rangle}=-\int^{T}_{0}\frac{dt}{T}F(r,v)v\,,$ (5.4)
$\Big{\langle}\frac{dL}{dt}\Big{\rangle}=-\sqrt{M\,a(e^{2}-1)}\int^{T}_{0}\frac{dt}{T}\frac{F(r,v)}{v}\,.$
Given an $F(r,v)$ we can compute $\Big{\langle}\frac{dE}{dt}\Big{\rangle}$ and $\Big{\langle}\frac{dL}{dt}\Big{\rangle}$. In our context, the dissipative effects that we will explore are the effects of dynamical friction with DM spike and the total orbital and angular momentum loss over the orbital timescale is given by
$\frac{dE_{\text{orb.}}}{dt}=\Big{\langle}\frac{dE_{\text{GW}}}{dt}\Big{\rangle}+\Big{\langle}\frac{dE_{\text{DM}}}{dt}\Big{\rangle}\,,$ (5.5)
$\frac{dL_{\text{orb.}}}{dt}=\Big{\langle}\frac{dL_{\text{GW}}}{dt}\Big{\rangle}+\Big{\langle}\frac{dL_{\text{DM}}}{dt}\Big{\rangle}\,.$
To obtain the change in orbital parameters we take resort to the following equations,
$E_{\text{orb.}}=\frac{\mu M}{2a}\Rightarrow\frac{da}{dt}=\frac{dE_{\text{orb.}}}{dt}/\frac{\partial E}{\partial a}\,,$ (5.6)
$\frac{de}{dt}=-\frac{e^{2}-1}{e}\Big{(}\frac{dE_{\text{orb.}}}{dt}/E_{\text{orb.}}+2\frac{dL_{\text{orb.}}}{dt}/L_{\text{orb.}}\Big{)}\,.$
Given these equations and a particular model for the dissipative force (which is the dynamical friction force), we can calculate these evolutions. For our context, the dissipative force inspired is of the form: $F_{0}r^{\gamma}v^{\delta}$ [114]. Here, $F_{0}=4\pi m_{2}^{2}r_{sp}^{\alpha}\log\Lambda.$ This follows from (3.4). Upon using such a form for the dissipative force we can calculate $\dot{a}$ by
[…]
$\dot{e}=-\frac{F_{0}}{2\pi}\frac{1}{e}\Big{[}(e^{2}-1)^{k_{1}+1/2}\frac{M^{\frac{\delta-2}{2}}}{\mu}a^{\tilde{k}_{1}}\int^{2\phi_{0}}_{0}d\phi(1+e\cos\phi)^{-(2+\gamma)}(1+2e\cos\phi+e^{2})^{\frac{\delta+1}{2}}$
$-(e^{2}-1)^{k^{\prime}_{1}+1/2}M^{\frac{\delta-3}{2}}a^{k^{\prime\prime}_{1}}\int^{2\phi_{0}}_{0}d\phi(1+e\cos\phi)^{-(2+\gamma)}(1+2e\cos\phi+e^{2})^{\frac{\delta-1}{2}}\Big{]}\,.$
To proceed further with these expressions, we expand these integrals around $e=1$ to get an analytic grasp of these integrals (we leave the general case for future study). We do this by taking $e=1+\epsilon$ and then expanding in terms of $\epsilon$. We also consider a static profile, leaving the study of effect of evolution of the dar matter profile to braking index for future. Concentrating on the form of $F(r,v)$ for the dynamical friction force given in (3.4), we can infer that $\gamma=-\alpha=-2.25$ (as defined below (3.1) [^8] We consider only one value of $\alpha$ as at this point we are only interested in getting an order of magnitude estimate for the braking index. However, our analysis is valid for other values of $\alpha$ mentioned below (3.1). and $\delta=-2$ (this corresponds to setting $\xi(v)=1$ in (3.4).). Using these values and also expanding in $\epsilon\,,$ we get
$\dot{a}=-\frac{F_{0}}{2^{1/4}\pi\mu}\frac{a^{1/4}}{\epsilon^{1/4}\,M^{2}}\Big{(}1-\frac{\epsilon}{8}\Big{)}(6.23634-1.68719\sqrt{\epsilon})\,,$ (5.8)
$\dot{\epsilon}=-\frac{F_{0}}{2\pi}\frac{2^{3/4}\epsilon^{3/4}}{a^{3/4}}\Big{[}\frac{1}{\mu\,M^{2}}(6.23634-1.68719\sqrt{\epsilon})-\frac{1}{M^{5/2}}(0.51965-0.420448\sqrt{\epsilon})\Big{]}\,.$
As is evident from the definition of the braking index, $n_{b}$, given in (5.3), the above expressions are enough to infer the braking index for our scenario.
Figure 5: Plots showing the evolution of eccentricity and the evolution of the braking index for an initial eccentricity $e_{0}=1.01,$ initial semi-major axis $a_{0}=100$ and mass ratio $q=10^{-5}\,.$ Inside the figure in the right panel, we have included an inset showing zooming on the fall-off part of the braking index. The spiky behaviour is due to the limitation of our numerical precision.
Now, we plot the braking index as a function of time. Fig. (5) shows some interesting features worth mentioning. Early in the evolution, the braking index is constant with the saturation value being $\sim 1.666$, and this is the value of the index in “vacuum”, i.e., when the environmental effects or the spike effects are not there  [^9] The plot for the braking index that we show here shows a sudden spike when the value for the index starts to decrease on the timescale of $\sim 10^{4}$ yr. We emphasize that this “spike” is due to some numerical artefact and not due to any other non-trivial physics.. As they start to come closer together and enter the halo region, the braking index shows a fall, with an estimated time for the fall around $\sim 10^{4}$ yr. Firstly, the plot for eccentricity evolution serves as a consistency check that the $e$ remains in the vicinity of unity as assumed while deriving the relevant equations in (LABEL:eq:5.8). Secondly, it allows us to explore the effects of the environment at large times, which will lead to larger dephasing. The timescale for such effects to be predominant is roughly $\sim 10^{4}$ yr, nearly identical to the braking index plot.
The observability of the second derivative of frequency evolution $\mathcal{F}$ has been an interesting problem to deal with like the one reported in [166] and so is the braking index for inspiraling binaries [167]. These results have an indication that such effects might be possible to detect if the dissipative forces are strong enough.
@@ 6.1 Effect of gravitational potential due to the Dark Matter Minispike
The above equation (6.6) tells us the perturbing force spans the plane in which the binaries are moving and the components contributing to the perturbing force can be easily read off by comparing with (6.2) and identifying $\hat{n}$ with $\hat{r}.$ Then from (LABEL:fig:5.5) and (LABEL:fig:5.6) we get,
@@ 6.2 Dynamical friction and accretion effects
Given a binary moving through a medium (which for us is a dark matter medium) it will experience a drag or dynamical friction force. Having discussed the effects of dark matter as perturbing force, we can now focus on the effects of such matter medium on the overall dynamics in such hyperbolic tracks. To this end as we discussed in Section 3 the “frictional” force of the medium can be modeled after a force law which takes the form -
@@
Once we have these components, we can plug them in the osculating equations (LABEL:eq:_5.3) and (LABEL:fig:5.5) and find the change in the orbital parameters:
@@ 6.4 Results: Numerical Analysis and Inferences
Now that we have all the relevant equations for studying the binary dynamics, specifically the effects of these various forces due to the presence of dark matter on the rate of change of the eccentricity. Before discussing our result, please note that we have used the fourth-order Runge-Kutta method to solve the coupled differential equations for each case separately, as mentioned in (LABEL:eq:5.7), (LABEL:eq:5.10) and (LABEL:eq:5.13). For each case, three coupled differential equations govern the evolution of eccentricity $e$, impact parameter $b$ and angle of closest approach $\phi_{0}$ w.r.t the true anomaly $\tilde{\phi}.$ Since we have taken true anomaly as our independent variable instead of time, all the plots are with w.r.t $\tilde{\phi}.$ The range of the true anomaly is from $\tilde{\phi}=-\phi_{0}$ to $\tilde{\phi}=\phi_{0}$ indicating the fact that the second object comes from infinity and means binaries after the encounter flies to infinity again. For the net effect, we first add the right-hand side of the (LABEL:eq:5.7), (LABEL:eq:5.10) and (LABEL:eq:5.13) and then sole the resulting coupled differential equation using the fourth-order Runge-Kutta method.
@@
6.5 The “curious” case of baryons
In most cosmological models, numerical simulations regarding the formation of structures show an inner core of the dark matter halo, which survives and gives way to more intricate structures like “subhalos” within their hosts. These subhalos boost gamma-ray production from dark matter annihilation and have also been known to enhance local cosmic ray production. However, all these estimates regarding such boosts and enhancements assume that baryon’s gravitational effects on the dark matter substructure are negligible. However, studies in [103, 104] have been able to show that there are viable feedback effects that can significantly alter distributions. In [105] show that given a dwarf galaxy setup, there could be dark matter heating due to baryons’ adiabatic compression of dark matter. If $\rho$ is the dark matter density and $v_{d}$ is the velocity dispersion of the gas particles, we can define $F=\frac{\rho}{v_{d}^{3}}$ as the phase-space density. Unlike the $\rho$ for the dark matter density and $v_{d}$, this phase space density is unaffected in dark matter simulations. However, hydrodynamic simulations give a different picture for $F$, the trend being a gradual decrease with time due to stellar feedback. This decrease is on a factor of 10 as compared to dark matter simulations. The stellar feedback only in the galaxy’s central region strongly affects the dark matter density. This is the region where the enclosed gas mass occasionally dominates that of the dark matter and is where the feedback most strongly affects the gas. At the end of the hydrodynamic simulations, the dark matter density at the smallest resolved radius becomes a factor of seven smaller than in the dark-matter-only simulations.
Given the particular shape of the halo profile, novel mechanisms like “resonant heating” can introduce changes to the cusp profile. Dark matter simulations suggest such cusp profiles, consistent with predictions from the standard model. However, “resonant heating” due to stellar feedback in hydrodynamic simulations turns these “cuspy” profiles into a “flat” core. One of the effects of this flattening is to reduce the efficiency of dynamical friction in the central regions. Again, this flattening is due to the stellar feedback (primarily arising out of baryons(stars and gas)). Such feedback may induce some heating in the globular cluster systems until stars stop forming. The velocity distribution is isotropic within the core and shows some radial anisotropy outside, which is not observed if massive gas clouds drive the mechanism. In that case, there are tangential anisotropies in addition to radial ones.
Given these effects, the response of dark matter distributions to baryons is topsy-turvy. While the above studies have shown a flattened curve to the profile, there are scenarios where the effects are otherwise. To resolve these discrepancies, theoretical arguments and simulations have proposed baryonic processes that can produce an expansion of the dark matter halo. Gas bulk motions, possibly supernova-induced in regions of high star formation activity, and the subsequent energy loss of gas clouds due to dynamical friction can transfer energy to the central dark matter component. In [106], it has been pointed out that there is another (possibly more relevant) effect, namely the gas bulk motion can induce substantial gravitational potential fluctuations and a subsequent reduction in the central dark matter density. In [107], it was shown that supernovae can eject matter from halos up to 100 km/s, but it has yet to be seen what impact this might have on dark matter profiles, nor how the addition of radiation pressure feedback might change things. So, while observations show evidence for flattened dark matter density profiles up to L* galaxies, the question remains whether there is enough energy input from baryons in more massive objects for these processes to be effective in altering the dark matter density profile of spiral galaxies with a dark matter mass of the order of $10^{11}-10^{12}$.
In light of these lessons, we focus on the density distribution and their related effects in our case, speaking of which we slightly change our profile following [108]. The new profile looks like the following:
$\rho_{DM}=\begin{cases}0&\text{when }r<r_{min}\,,\\ \rho_{sp}\Big{(}\frac{r_{sp}}{r}\Big{)}^{\alpha}&\text{when }r_{min}\leq r<r_{sp}\,,\\ \rho^{\text{cored}}_{\text{DM-HALO}}&\text{when }r\geq r_{sp}\end{cases}$ (6.15)
where the cored halo profile has the following structure:
$\rho^{\text{cored}}_{\text{DM-HALO}}=\begin{cases}\rho^{\text{NFW}}_{\text{halo}}(r_{c})\Big{(}\frac{r}{r_{c}}\Big{)}^{-\gamma_{c}}&\text{when }r<r_{c}\,,\\ \rho^{\text{NFW}}_{\text{halo}}&\text{when }r\geq r_{c}\end{cases}$ (6.16)
where $\rho^{\text{NFW}}_{\text{halo}}=\rho_{s}(\frac{r}{r_{s}})^{-1}(1+\frac{r}{r_{s}})^{-2}$ is the ubiquitous NFW profile with $r_{s}=18.6$ kpc and $\rho_{s}=\rho_{\odot}(R_{\odot}/r_{s})(1+R_{\odot}/r_{s})^{2}$, where $R_{\odot}=8.2$ kpc is the Sun position and $\rho_{\odot}=0.383$ GeV/$\text{cm}^{3}$, being the local DM density. Upon “correcting” the profile taking into the baryonic effects as discussed before into consideration, we observe the following features:
Figure 13: The above figures show the variation of eccentricity as a function of true anomaly parameter for different values of impact parameter $b_{0}$. The blue one is the one with the presence of spike only and the orange one is the one with an additional halo region covering it. The set of values we take are: $M_{\text{DM}}=10^{6}M_{\odot},M_{\text{BH}}=10^{3}M_{\odot},r_{\text{s}}=23.1\text{pc},\rho_{\text{s}}=3.9\times 10^{-19}\text{kg}/\text{m}^{3},r_{\text{sp}}=0.54\text{pc},\rho_{\text{sp}}=226M_{\odot}/\text{pc}^{3}$ and $r_{\text{min}}=\frac{6GM_{\text{BH}}}{c^{2}}\,.$
  • The eccentricity of the orbit shows a sharper fall-off than the one without the halo, as can be seen in Fig. (13). This can be understood logically since the spike and halo now dampen the hyperbolic binaries, leading to a greater fall for the eccentricities.
    Figure 14: Plots showing the behavior of eccentricity as a function of impact parameter for different values of $\alpha$, the dark matter parameter index. The blue one is only for the spike profile and the orange one is the one with an additional halo region covering it. The set of values we take are: $M_{\text{DM}}=10^{6}M_{\odot},M_{\text{BH}}=10^{3}M_{\odot},r_{\text{s}}=23.1\text{pc},\rho_{\text{s}}=3.9\times 10^{-19}\text{kg}/\text{m}^{3},r_{\text{sp}}=0.54\text{pc},\rho_{\text{sp}}=226M_{\odot}/\text{pc}^{3}$ and $r_{\text{min}}=\frac{6GM_{\text{BH}}}{c^{2}}\,.$
  • Another interesting feature we observe in our profile is when we vary the dark matter parameter $\alpha$. As we increase the index - the variation being in the permissible range of 2.25 to 2.5, we see the two profiles merging. This is evident from Fig. (14). To remind the readers again, these two profiles are the dark matter spike with the halo and the one without the halo.
6.5.1 Including the “annihilation” region
We can also do better and include some other effects too. As can be rightly inferred, the spike radius saturates due to DM annihilations. One can look into the physics of such annihilation in our orbital parameter behaviour. These annihilations happen in a region which is the innermost of this halo plus spike domain, weakening the density profile there. When considering models like the WIMP ones, the annihilation cross section $\langle\sigma v\rangle$ is constant, flatting the dark matter profile. Let $r=r_{\text{anni}}$ be this “plateau” region and $\rho_{\text{anni}}$ being the corresponding density, then [99]
$\rho_{\text{anni}}=\frac{m_{\xi}}{\langle\sigma v\rangle T}$ (6.17)
where $m_{\xi}$ is the dark matter particle mass and $T$ is the galaxy age.
A possible explanation for these plateaus was given in [98], due to dark matter (DM) particles moving in strictly circular orbits around the central black hole. In [98], it was shown that if the dark matter distribution is isotropic (which was likely so in all the simulations and analysis), the density shows a steep rise and eventually forms a cusp and not a plateau. The density varies as $\sim r^{-1/2}$ and the cusp is maintained as the particles outside $r=r_{\text{anni}}$ continue to contribute to the density inside $r=r_{\text{anni}}$. The distinction between a plateau and a cusp has important observational consequences  [^13] For a brief review see [100]. and hence an analysis including such subtle but non-trivial physics is somewhat lucrative.
$\rho_{DM}=\begin{cases}0&\text{when }r<2r_{\text{S}}\,,\\ \rho_{\text{anni}}\Big{(}\frac{r}{r_{\text{anni}}}\Big{)}^{-0.5}&\text{when }2r_{\text{S}}\leq r<r_{\text{anni}}\,,\\ \rho_{sp}\Big{(}\frac{r_{sp}}{r}\Big{)}^{\alpha}&\text{when }r_{\text{anni}}\leq r<r_{sp}\,,\\ \rho^{\text{cored}}_{\text{DM-HALO}}&\text{when }r\geq r_{sp}\,.\end{cases}$ (6.18)
Upon including these effects, the behaviour in eccentricity shows the usual trend as demonstrated previously, meaning that it shows a sharper fall-off with the changes of $\mathcal{O}(10^{-2})$ than the one without them. Also, as seen in the case of halo, the profiles show a typical merging behaviour at the high values of $\alpha$. This is shown in Fig. (15).
Figure 15: Variation of eccentricity upon the addition of the “annihilation” region (green curve). We take $r_{anni}=3.1\times 10^{-3}\text{pc}$ and $\rho_{\text{anni}}=1.7\times 10^{8}M_{\odot}\text{pc}^{-3}$. The set of values we take are: $M_{\text{DM}}=10^{6}M_{\odot},M_{\text{BH}}=10^{3}M_{\odot},r_{\text{s}}=23.1\text{pc},\rho_{\text{s}}=3.9\times 10^{-19}\text{kg}/\text{m}^{3},r_{\text{sp}}=0.54\text{pc},\rho_{\text{sp}}=226M_{\odot}/\text{pc}^{3}$ and $r_{\text{min}}=\frac{6GM_{\text{BH}}}{c^{2}}\,.$